OpenStudy (anonymous):

m^4 * 2n^-3 is it 3m

OpenStudy (anonymous):

U can't simplify that any more...

OpenStudy (anonymous):

Where do u get 3m?

jimthompson5910 (jim_thompson5910):

\[\large m^4*2n^{-3}=\frac{2m^4}{n^3}\]

OpenStudy (anonymous):

D: ugh i hate math with a burning passion

OpenStudy (anonymous):

U don't love math, u just get used to it..:-)

OpenStudy (anonymous):

did i really put 2n? fml i meant 2m^-3

OpenStudy (anonymous):

I thought that....

OpenStudy (anonymous):

Didn't know which, though...

OpenStudy (anonymous):

so sorry. `~`

OpenStudy (anonymous):

Now it's like the other one, so 2 and add the exponents...

OpenStudy (anonymous):

so was i right? 3m?

jimthompson5910 (jim_thompson5910):

Add the exponents to get 4+(-3) = 1 So the final exponent is 1, which means that the answer is 2m^1 or just 2m

OpenStudy (anonymous):

wouldn't you have to add m + 2m? :l

OpenStudy (anonymous):

Look, m*m*m = m^3

OpenStudy (anonymous):

So the m is already in the exponent.

OpenStudy (anonymous):

No need to put again.

jimthompson5910 (jim_thompson5910):

For example, x^2 times x^3 is x^5 (since 2+3=5) or x^2 = x*x and x^3 = x*x*x so (x^2)*(x^3) = (x*x)*(x*x*x) = x*x*x*x*x = x^5

OpenStudy (anonymous):

Only the 2 lives by itself.

OpenStudy (anonymous):

so the final answer is 2m?

jimthompson5910 (jim_thompson5910):

yes, the final answer is 2m

OpenStudy (anonymous):

I give u one 3^3* 5n^-2* 4n^4 = ?

OpenStudy (anonymous):

See if u can figure out...

OpenStudy (anonymous):

would it be 3^3 * 20n^2 or 60n?

OpenStudy (anonymous):

This 3^3 * 20n^2 is right.....

OpenStudy (anonymous):

Remember 3^3 = 3*3*3

OpenStudy (anonymous):

woot

OpenStudy (anonymous):

Same like n^3... = n*n*n so does 3^3 = 3*3*3 (letter instead of number)

OpenStudy (anonymous):

i understand

OpenStudy (anonymous):

i'll be posting a new question here shortly :)

OpenStudy (anonymous):

So 3^3 * 20n^2 = 27* 20n^2 = 540n^2

OpenStudy (anonymous):

\[2x ^{3}y ^{-3} * 2x ^{-1}y ^{3}\]

OpenStudy (anonymous):

I think u can do it...

OpenStudy (anonymous):

= 4x^2

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

:D thanks so much for all your help!

OpenStudy (anonymous):

ur welcome, anytime...